OpenStudy (anonymous):

Find the center of mass of a solid cylinder C that has density f=3z^2, bounded by x^2+y^2<(or equal)1 and 0<(or equal) z <(or equal)1

OpenStudy (vincent-lyon.fr):

General definition is : \(\vec {OG} = \LARGE\frac 1m \Large\iiint \limits_ {P\in Cyl.} \,\normalsize \vec{OP}\;\mathrm{d}m\) which leads here to : \(h=\LARGE\frac 1m \Large \int _0^1 \,\normalsize z\;\mathrm{d}m\) where \(m= \Large \int _0^1 \,\normalsize \mathrm{d}m\)